### v1

# EE-MLLM: A Data-Efficient and Compute-Efficient Multimodal Large Language Model

Abstract
In the realm of multimodal research, numerous studies leverage substantial image-text pairs to conduct modal alignment learning, transforming Large Language Models (LLMs) into Multimodal LLMs and excelling in a variety of visual-language tasks. The prevailing methodologies primarily fall into two categories: self-attention-based and cross-attention-based methods. While self-attention-based methods offer superior data efficiency due to their simple MLP architecture, they often suffer from lower computational efficiency due to concatenating visual and textual tokens as input for LLM. Conversely, cross-attention-based methods, although less data-efficient due to additional learnable parameters, exhibit higher computational efficiency by avoiding long sequence input for LLM. To address these trade-offs, we introduce the Data-Efficient and Compute-Efficient Multimodal Large Language Model (EE-MLLM). Without introducing additional modules or learnable parameters, EE-MLLM achieves both data and compute efficiency. Specifically, we modify the original self-attention mechanism in MLLM to a composite attention mechanism. This mechanism has two key characteristics: 1) Eliminating the computational overhead of self-attention within visual tokens to achieve compute efficiency, and 2) Reusing the weights on each layer of LLM to facilitate effective modality alignment between vision and language for data efficiency. Experimental results demonstrate the effectiveness of EE-MLLM across a range of benchmarks, including general-purpose datasets like MMBench and SeedBench, as well as fine-grained tasks such as TextVQA and DocVQA.
Figure 1: Architecture comparisons between self-attention-based method, cross-attention-based method and our EE-MLLM. (1) The self-attention-based mechanism utilizes a projector to align visual tokens with text tokens, subsequently concatenating these tokens as input for the LLM. (2) The cross-attention-based mechanism integrates additional cross-attention blocks into the decoder layers. (3) EE-MLLM introduces a composite attention mechanism to eliminate the computational overhead of self-attention within visual tokens and reuse the weights as aligners on each layer of LLM to facilitate modality alignment.
Introduction
Recently, Large Language Models (LLMs) have gained significant attention due to their surprising performance on various natural language tasks. However, real-world scenarios often involve more than just linguistic modality, making it crucial to extend LLM to Multimodal LLMs. The key to expansion is to perform modality alignment, that is, to learn to map the remaining modalities with the same semantics to their corresponding language modality in the feature space of the pre-trained LLM.
In the field of multimodal LLM research involving vision and language, a series of research works on modality alignment have been carried out and achieved excellent performance. These methods can be roughly divided into two categories: self-attention-based methods and cross-attention-based methods.
The self-attention-based methods firstly leverage a simple MLP layer to project the visual tokens into the input space of LLM, and then concatenate the visual tokens with the textual tokens, which is shown in Fig.1 (1). These flattened concatenated long token sequences are fed into the LLM to further model the dependencies of various tokens. These methods exhibit great data efficiency due to two factors: 1) The alignment module is simple and has a small number of learnable parameters. 2) Visual tokens aligned in the input space of LLM will naturally lead to the alignment between vision and language in each layer of LLM. Therefore, the training procedure does not need to focus on the modality alignment within the LLM, enabling data efficiency. For example, LLaVA (Liu et al. 2023b) achieves promising performance by pre-training on only 558k data and fine-tuning on only 665k instruction data. However, the flattening operation of directly concatenating visual tokens and textual tokens expands the length of the token sequence from $T$ to $V+T$ (where $T$ is the number of textual tokens, $V$ is the number of visual tokens), and the computational cost of the language model layer (self-attention structure) increases from $O(T^{2})$ to $O((V+T)^{2})$, results in lower computer-efficiency. Especially when the input image is high-resolution, the number of visual tokens increases, and the compute inefficiency will prolong the training and inference time. Specifically, LLaVA (Liu et al. 2023b) and Deepseek-VL (Lu et al. 2024) have 576 visual tokens, Sphinx (Lin et al. 2023) utilizes 2,890 visual tokens.
The cross-attention-based methods directly insert new cross-attention layers between existing pre-trained language model layers. As shown in Fig.1 (2), the keys and values in these layers are obtained from the visual tokens while the queries are derived from the text tokens. These methods have higher compute efficiency because the length of the input sequence is always kept as $T$ and does not increase with the increase of visual tokens. However, these methods usually require a large amount of pre-training data, leading to data inefficiency. For example, Flamingo (Alayrac et al. 2022) is pre-trained on 1.8 billion image-text pairs (Jia et al. 2021) and interleaved data. EVLM (Chen et al. 2024) is pre-trained on 2.5 billion image-text pairs and 50 million interleaved image-text data. This data inefficiency arises from two factors: 1) The additional cross-attention layers will introduce a large number of learnable parameters, necessitating more training data. 2) The training procedure has to focus on the modality alignment between vision and language on different layers of LLM, posing a huge optimization complexity for modality alignment.
By reviewing existing methods, we conclude that the key points to ensure data efficiency and compute efficiency are 1) Avoiding an increase in the length of input sequence for LLM, 2) The modality alignment module should have fewer parameters, and the alignment needs to be effective. Taking two summary viewpoints into account, we present a data-Efficient and compute-Efficient Multimodal Large Language Model (EE-MLLM). Specifically, we propose a composite attention mechanism to replace the original self-attention mechanism in MLLM. The composite attention mechanism has two features: 1) Eliminating the computational overhead of self-attention within visual tokens to achieve compute efficiency. The original self-attention mechanism of MLLM consists of three components: self-attention within visual tokens, self-attention within text tokens, and cross-attention between text and visual tokens. We observe that the self-attention within visual tokens is redundant. This is because a) the interaction between visual tokens is already well-processed by the vision encoder, and b) LLM can facilitate the implicit interaction among visual tokens via the information aggregation property of LLM (Wang et al. 2023b). Therefore, we eliminate the self-attention within visual tokens and maintain the length of the input sequence for LLM as the length of text tokens, thereby facilitating computational efficiency. 2) Instead of introducing additional modules like cross-attention-based methods, we reuse the weights on each layer of LLM to facilitate effective modality alignment between vision and language for data efficiency. As a result, without introducing additional parameters, visual tokens aligned in the input space of LLM can naturally facilitate the alignment between vision and language in each layer of LLM. This means the training procedure does not need to focus on the modality alignment within the LLM, which enables data efficiency.
To demonstrate the effectiveness of EE-MLLM, we evaluate its performance on both general benchmarks and fine-grained benchmarks, such as MMBench-EN, MME, ScienceQA, AI2D, TextVQA, and DocVQA. In comparison with state-of-the-art MLLMs, our EE-MLLM exhibits promising results on both general and fine-grained benchmarks. Simultaneously, EE-MLLM proves to be computationally efficient in the inference stage. When applied to high-resolution image inputs, EE-MLLM maintains comparable performance at a significantly reduced computational cost. For instance, with a $980\times 980$ input image, the number of visual tokens can reach 4,900, and the FLOPs of EE-MLLM are $70\%$ of those in self-attention-based methods. We also assess the inference speed of EE-MLLM in a real-world scenario, confirming its efficiency. Specifically, when the number of generated tokens is set to 32, EE-MLLM achieves an inference speed of 77 tokens per second on a single NVIDIA H800 GPU, which is 1.9 times faster than self-attention-based methods at the same resolution $980\times 980$.
Figure 2: Our composite attention mechanism consists of the composite attention module and the aligner. For the aligner, visual tokens are aligned to the feature space of each layer of LLM by the aligner alone; for the composite attention module, the concatenation of visual and text tokens are used as keys and values, and the text tokens are used as queries for attention, thus eliminates the self-attention within visual tokens.
Our main contributions are summarized as follows:
  1. 1.
    We revisit modality alignment in multimodal LLMs, pointing out the issue of efficiency and efficacy in self-attention-based and cross-attention-based methods.
  2. 2.
    We propose a composite attention mechanism for EE-MLLM to enhance the data and computational efficiency.
  3. 3.
    Our EE-MLLM achieves outstanding performance on a variety of benchmarks, while the inference speed is much improved.
Related Work
Self-attention-based methods. Self-attention-based methods (Dai et al. 2024; Zhu et al. 2024; Chen et al. 2023a; Bai et al. 2023; Chen et al. 2023b; Fang et al. 2023) employ a projector to align visual features with the input space of LLM. LLaVA (Liu et al. 2023b, a) uses a simple MLP layer to connect the vision encoder of CLIP (Radford et al. 2021) with Vicuna (Zheng et al. 2023). DeepSeek-VL (Lu et al. 2024) employs hybrid vision encoders, including SigLIP (Zhai et al. 2023) for high-level semantic features and SAM-B (Kirillov et al. 2023) for low-level features. Considering the scenarios that require fine-grained information, such as document recognition and form understanding, recent work extends MLLM to support high-resolution inputs. LLaVA-Next (Liu et al. 2024) proposes dividing the image into small patches for extracting local features and concatenating the downsampled image feature as global context. Monkey (Li et al. 2024b) also partitions high-resolution images into small patches, processing each patch independently with a trainable resampler. Since the self-attention-based methods are compute-inefficiency, the overhead for high-resolution inputs can be more severe. Deco (Yao et al. 2024) and HoneyBee (Cha et al. 2024) carefully design projectors to reduce the number of visual tokens with minimal information loss. However, these methods still compromise fine-grained capabilities when reducing the number of visual tokens. Our EE-MLLM, a cross-attention-based architecture, achieves state-of-the-art performance while maintaining computational efficiency.
Cross-attention-based methods. Cross-attention-based methods (Laurençon et al. 2023; Alayrac et al. 2022; Awadalla et al. 2023; Chen et al. 2024) typically incorporate additional cross-attention modules into the decoder layers of LLM to integrate visual information without extending the length of the text token sequence. Flamingo (Alayrac et al. 2022) introduces a novel gated xattn-dense block, incorporating a tanh-gating mechanism for training stability. Flamingo is trained on large-scale interleaved image-text data and supports in-context few-shot learning. EVLM (Chen et al. 2024) enhances Flamingo by: (1) utilizing hierarchical ViT features without a resampler, (2) replacing the image token with a set of learnable tokens in the text input, and (3) introducing MoE to boost performance. Although these cross-attention-based methods are computationally efficient, the additional cross-attention blocks lead to a large number of trainable parameters and a heavy reliance on pre-training data. Our EE-MLLM avoids introducing any extra trainable parameters and repurposes the weights of LLM to map visual tokens into the LLM feature space, thus achieving promising performance with a small amount of training data.
Method
Model Architecture
EE-MLLM comprises a visual encoder, a two-layer MLP serving as the projector, and the LLM with composite decoder layers. We propose a composite attention mechanism and design a composite decoder layer for EE-MLLM to achieve data efficiency and compute efficiency. The composite decoder layer consists of an aligner and a composite attention module.
Composite Attention Module. The original self-attention mechanism comprises self-attention within visual tokens, self-attention within text tokens, and cross-attention between text tokens and visual tokens. We observe that self-attention within visual tokens is redundant for two reasons: 1) The interaction between visual tokens is already well-learned in the vision encoder, and 2) LLM can enable the implicit interaction between visual tokens via the information aggregation property (Wang et al. 2023b). Therefore, we eliminate self-attention within visual tokens and obtain the composite attention module. We assume that visual tokens $I\in\mathbb{R}^{k\times h}$ and text tokens $T\in\mathbb{R}^{n\times h}$, where $k$ and $n$ are the lengths of visual tokens and text tokens, respectively. And $h$ denotes the dimension of hidden states. The original self-attention can be formulated as follows:
$\textrm{Attention}(X^{q},X^{k},X^{v})=\textrm{softmax}(\frac{X^{q}{X^{k}}^{T}}{\sqrt{h}})X^{v},$ (1)
where $X^{q}=[I;T]W_{q}$, $X^{k}=[I;T]W_{k}$, $X^{v}=[I;T]W_{v}$, $W_{q}$, $W_{k}$ and $W_{v}$ are the weights of the self-attention module. The symbol [;][;][ ; ] denotes the concatenation operation of two variables in the sequence dimension. As illustrated in Fig. 2, our composite attention module uses text tokens as the queries and the concatenation of visual tokens and text tokens as keys and values for cross-attention. Specifically, $X^{q}=TW_{q}$, $X^{k}=[I;T]W_{k}$, and $X^{v}=[I;T]W_{v}$. The length of the output sequence will be $n$, which is the same as the length of text tokens $T$. The attention mask for self-attention is a lower triangular matrix, whereas the attention mask for the composite attention module is a trapezoidal matrix. This implies that text tokens will attend to all previous tokens, including visual tokens, while the interaction within visual tokens is eliminated.
Type Dataset Ratio
General Instruction Data ShareGPT4V 35%
LVIS-Instruct4V
LAION-GPT4V
TextOCR-GPT4V
Localized Narratives
IconQA
Text-Centric Data DVQA 8.5%
PlotQA
CLEVR
DocVQA
ChartQA
ScienceQA
Others Text-only data 47%
In-house data 9.5%
Table 1: Details of supervised fine-tuning data.
Aligner. We introduce the aligner, which leverages the existing weights at each layer of LLM to enhance modality alignment between vision and language without the need for additional modules such as cross-attention-based methods. By reusing these weights, visual tokens aligned in the input space of the LLM can naturally facilitate modality alignment at each layer. This approach eliminates the necessity for the training procedure to concentrate on modality alignment within the LLM, thereby promoting data efficiency. As depicted in Fig. 2, in each layer, text tokens are mapped by the value matrix ($W_{v}$), output matrix ($W_{o}$), and FFN. Hence, we make visual tokens follow the same mapping relationship in the aligner. Specifically, the aligner comprises $W_{v}$, $W_{o}$, and FFN, which share weights with the corresponding LLM modules. The residual connection is also implemented in the aligner. The aligner can be formulated as follows:
$H$ $=IW_{v}W_{o}+I$ (2)
$O$ $={\rm FFN}(H)+H.$ (3)
This can be viewed as a decoder block without attention.
Analysis of the Computational Overhead
Our EE-MLLM can reduce the computational overhead without introducing additional parameters. In this section, we analyze the compute efficiency of EE-MLLM from the perspective of FLOPs, comparing it with LLaVA. The LLM consists of $d$ blocks, with a hidden states dimension of $h$, an input text length of $T$, and $V$ visual tokens. Considering an LLM, we assume the input sequence length is $L$, the FLOPs of the self-attention module and FFN are $8Lh^{2}+4L^{2}h$ and $16Lh^{2}$. For self-attention-based methods, such as LLaVA, the length of the input sequence is $L=V+T$. Therefore, the FLOPs of LLaVA are $24(T+V)dh^{2}+4(T+V)^{2}dh$. For EE-MLLM, the FLOPs of composite attention modules are $(6T+2V)h^{2}+4VTh+4T^{2}h$. Given that we introduce aligners, the FLOPs of FFN are the same as those of LLaVA, which are $16(V+T)h^{2}$. The aligners introduce an additional matrix multiplication by $O$, adding $2Vh^{2}$ FLOPs. Therefore, the total FLOPs of EE-MLLM are $2(11T+10V)dh^{2}+4VTdh+4T^{2}dh$. Particularly, when the resolution of the input image is $980\times 980$, the number of visual tokens is 4,900, and the length of text tokens is 256, which is a typical length during training, the FLOPs of EE-MLLM amount to 70% of those of LLaVA.
Experiments
Implementation Details
Model Configurations. We employ Vicuna-7b-v1.5 (Zheng et al. 2023) as our LLM and SigLIP (Zhai et al. 2023) as the vision encoder. Specifically, SigLIP is initialized from Idefics2 (Laurençon et al. 2024), which supports dynamic resolutions up to 980 $\times$ 980. The projector consists of a two-layer MLP, identical to that of LLaVA (Liu et al. 2023b).
Model LLM MMB-T MME ScienceQA HallB MMMU CCBench SeedBench BLINK
InstructBLIP Vicuna-7B 36.0 1137.1 54.7 31.2 30.6 12.7 44.5 -
MiniGPT-4-v1 Vicuna-7B 12.2 770.6 39.0 31.9 23.6 1.8 31.6 -
MiniGPT-4-v2 Llama2-7B 24.3 708.4 54.1 30.0 25.0 1.4 29.4 -
Idefics-instruct Llama-7B 48.2 942 51.6 27.3 18.4 7.8 45.0 38.3
OpenFlamingo v2 MPT-7B 6.6 535 45.7 29.4 28.2 6.3 28.8 -
Qwen-VL Qwen-7B 38.2 334.1 57.7 29.9 29.6 6.1 52.5 27.9
Qwen-VL-Chat Qwen-7B 60.6 1467.8 65.5 36.8 37.0 41.2 64.8 28.2
VLoRA Vicuna-7B 63.4 1311.3 66.4 26.4 36.0 28.6 - -
ShareGPT4V Vicuna-7B 64.6 1561.4 68.2 28.6 37.2 30.8 69.3 40.9
LLaVA-v1.5 Vicuna-7B 62.3 1510.7 66.8 27.6 35.7 27.5 65.8 39.7
LLaVA-v1.6 Vicuna-7B 66.5 1475.6 68.5 27.6 37.6 24.3 69.6 41.6
EE-MLLM Vicuna-7B 70.4 1528.1 77.7 38.6 33.4 37.3 70.2 43.2
Table 2: Comparsion with state-of-the-art methods on general benchmarks, including MMBench, MME, ScienceQA, HallusionBench, MMMU, CCBench, SeedBench, and BLINK. MMB-T indicates we report the results on MMBench-TEST-EN-V11.
Pre-training Configurations. During the pre-training stage, we follow LLaVA to adopt blip-558k (Li et al. 2022) for training one epoch. We freeze the LLM and the vision encoder, making only the projector trainable. We adopt AdamW (Loshchilov and Hutter 2018) optimizer. The learning rate is 1e-3, followed by a linear warm-up scheduler and then a cosine decay scheduler. The global batch size is 256.
Fine-tuning Configurations. During the fine-tuning stage, we freeze the vision encoder and update the weights of LLM and projector. The total quantity of supervised fine-tuning data amounts to 3 million. As displayed in Tab. 1, following Deepseek-VL (Lu et al. 2024), our supervised fine-tuning data incorporates open-source gpt4v datasets, including ShareGPT4V (Chen et al. 2023b), LVIS-Instruct4V (Wang et al. 2023a), LAION-GPT4V (LAION 2023), and TextOCR-GPT4V (Carter 2024). Moreover, to enhance the fine-grained capabilities of MLLM, we introduce DVQA (Kafle et al. 2018), PlotQA (Methani et al. 2020), CLEVR (Johnson et al. 2017), DocVQA (Mathew, Karatzas, and Jawahar 2021), ChartQA (Masry et al. 2022), and ScienceQA (Lu et al. 2022) for fine-tuning. To preserve language abilities during fine-tuning, we utilize a large volume of text-only data from various sources. The text-only data encompasses various dataset types, such as mathematical reasoning, multi-round dialogues, logic comprehension, and code. Meanwhile, the in-house data primarily consists of OCR-related, table and diagram comprehension, and mathematical reasoning data. The optimizer is AdamW with a global batch size of 128. The learning rate is 2e-5 and the learning rate scheduler is the same as the pre-training stage. The training is conducted on 16 NVIDIA H800 GPUs.
Evalution Benchmarks
We conduct our evaluations using the VLMEvalKit (Duan et al. 2024), and the results of other state-of-the-art models are obtained from the same source.
General Benchmarks. 1) MMBench-EN (Liu et al. 2023c) is a comprehensive multimodal benchmark specifically designed to assess the performance of MLLMs. It comprises over 3,000 multiple-choice questions spanning 20 ability categories. We evaluate EE-MLLM on the MMBench-EN-V1.1. 2) MME (Fu et al. 2023) evaluates advanced MLLMs in terms of perception and cognition, encompassing a total of 14 subtasks. To minimize the impact of prompt engineering on MLLMs, MME’s instructions are designed to elicit simple binary responses, such as ”please answer yes or no.”. We report the results of the perception split of MME. 3) ScienceQA (Lu et al. 2022)is derived from elementary and high school science curricula. The questions in ScienceQA cover three subjects: natural science, language science, and social science. 4) HallusionBench (Guan et al. 2023) is designed for evaluating image-context reasoning, comprising 346 images paired with 1,129 questions crafted by human experts. HallusionBench takes into account both language hallucinations and visual illusions across a diverse range of topics. 5) MMMU (Yue et al. 2023) collects 11.5K multimodal questions from college exams, quizzes, and textbooks, covering six core disciplines, spanning 30 subjects and 183 subfields, including 30 heterogeneous image types. 6) CCBench (Liu et al. 2023c), developed by the MMBench team, is specifically designed to evaluate MLLMs in the domain of Chinese Culture. 7) SeedBench (Li et al. 2023) encompasses 19K multiple choice questions, covering 12 evaluation dimensions, including both image and video. We only use questions with images for evaluation. 8) BLINK (Fu et al. 2024) contains 14 visual perception tasks which pose significant challenges for current multimodal LLMs.
Model LLM AI2D OCRBench TextVQA ChartQA DocVQA Seed2 Plus
InstructBLIP Vicuna-7B 40.6 27.6 33.6 10.9 - 29.5
MiniGPT-4-v1 Vicuna-7B 28.4 17.2 - - - 15.2
MiniGPT-4-v2 Llama2-7B 30.5 3.1 - - - 23.3
Idefics-instruct Llama-7B 42.2 25.2 - - - 35.4
OpenFlamingo v2 MPT-7B 31.7 14.9 16.3 - - 28.7
Qwen-VL Qwen-7B 57.7 12.7 63.1 59.0 - 40.1
Qwen-VL-Chat Qwen-7B 63.0 48.8 60.7 49.8 - 40.6
ShareGPT4V Vicuna-7B 58.0 37.1 51.1 21.3 - 46.1
LLaVA-v1.5 Vicuna-7B 55.5 31.8 45.5 17.8 - 41.3
LLaVA-v1.6 Vicuna-7B 67.0 53.2 64.4 55.4 - 51.6
EE-MLLM Vicuna-7B 65.7 57.4 69.0 68.4 67.4 53.5
Table 3: Comparsion with state-of-the-art methods on fine-grained benchmarks, including AI2D, OCRBench, TextVQA, ChartQA, DocVQA, InfoVQA, and SeedBench-2 Plus.
Model MMB-VAL SeedBench ScienceQA BLINK AI2D OCRBench TextVQA ChartQA Avg.
$336\times 336$
LLaVA-v1.5 67.3 65.1 67.0 40.7 56.4 24.4 46.1 17.9 48.1
EE-MLLM 66.8 64.1 66.7 38.4 56.2 23.6 43.3 17.3 47.1
$672\times 672$
LLaVA-v1.5 70.1 68.0 69.4 40.1 58.3 39.8 61.7 27.9 54.4
EE-MLLM 68.2 67.3 69.1 40.3 57.5 37.0 60.1 25.1 53.1
Table 4: Comparisons with LLaVA under the same settings, including LLM, vision encoder and training data.
Fine-grained Benchmarks. 1) AI2D (Kembhavi et al. 2016) emphasizes diagram interpretation and reasoning, comprising 5,000 diagrams and 15,000 questions and answers. 2) OCRBench (Liu et al. 2023d) aims to facilitate the assessment of MLLM OCR capabilities, including 29 datasets. 3) TextVQA (Singh et al. 2019) consists of 45,336 questions and 28,408 images that necessitate reasoning about text for answering. We use the validation set, containing 5,000 images, for evaluation. 4) ChartQA (Masry et al. 2022) is a large-scale benchmark featuring 20,882 charts, with questions focusing on logical and visual reasoning. 5) DocVQA (Mathew, Karatzas, and Jawahar 2021) concentrates on document image understanding, encompassing 50,000 questions and over 12,000 images. We evaluate using the validation set, which includes 5,349 questions and 1,286 images. 6) Seed2 Plus (Li et al. 2024a) specifically designed for text-rich visual comprehension evaluation of MLLMs, includes 2.3K multiple-choice questions covering charts, maps, and webs.
Figure 3: The visualization results of examples sampled from BLINK and RealWorldQA.
Comparisons with State-of-the-art Models
General Benchmarks. In Tab. 2, we compare EE-MLLM with various state-of-the-art MLLMs (Liu et al. 2024; Bai et al. 2023; Ma et al. 2024), across eight general benchmarks. These benchmarks evaluate the comprehensive ability of MLLMs, including understanding and perception, as well as the severity of illusion. These factors collectively reflect the generalization ability and applicability of MLLMs in real-world scenarios. EE-MLLM achieves comparable performance with state-of-the-art MLLMs in general benchmarks. Specifically, EE-MLLM achieves a score of 70.4 in MMBench and 1528.1 in MME, scores that are notably higer compared to LLaVA-v1.6, which also support high-resolution images input. This suggests that EE-MLLM possesses comprehensive perception and reasoning capabilities. What’s more, EE-MLLM also achieves promising performance on CCBench and SeedBench.
Figure 4: Illustration of Inference Speed Ratio between EE-MLLM and LLaVA for input resolutions of $980\times 980$, considering numbers of generated tokens ranging from 2 to 256, for single and double image inputs.
Fine-grained Benchmarks. In Tab. 3, we perform evaluations on seven fine-grained benchmarks. These benchmarks necessitate impressive visual perception capability of MLLMs, as they require the exploration of fine-grained information within images to answer the questions. For traditional VQA benchmarks, such as TextVQA, ChartQA, and DocVQA, MLLM achieves a very impressive performance, 4.6 higher on TextVQA and 13.0 higher on ChartQA compared to LLaVA-v1.6, which also supports high resolution. In the OCRBench, specifically designed to assess the OCR capabilities of MLLM, EE-MLLM outperforms by 4.2 compared to LLaVA-v1.6. These results demonstrate that despite EE-MLLM significantly reducing the computational overhead associated with visual tokens, it effectively maintains the model’s fine-grained capabilities.
Comparisons of Inference Speed
Despite the significant reduction in FLOPs demonstrated by EE-MLLM, it is crucial to consider real-world scenarios where advanced techniques such as KVCache and Batch Inference are typically employed during deployment. Hence, conducting comparisons of inference speed with these techniques becomes necessary. We conduct the comparisons of inference speed on a single NVIDIA H800. The resolution of the input image is set to $980\times 980$, and the number of generated tokens varies from 2 to 256. We illustrate the speed ratio between EE-MLLM and LLaVA in Fig. 4. Our findings show that when generating 8 tokens, EE-MLLM’s inference speed is three times that of LLaVA. However, as the number of generated tokens increases, the speed ratio decreases. When generating 64 tokens, EE-MLLM’s inference speed is 1.6 times that of LLaVA. The reason for this observation is that our EE-MLLM primarily reduces the computational cost during the prefilling stage, which calculates the KV cache of visual tokens. The generation of the first token is faster than self-attention-based methods like LLaVA. However, the advantage in inference speed diminishes after the first token. Specifically, with two input images, the inference speed of EE-MLLM is nearly four times that of LLaVA. This clearly indicates that EE-MLLM is significantly more efficient in multi-image input scenarios, including interleaved image-text dialogues and multimodal in-context learning.
Stage EE-MLLM LLaVA
Pre-training 32G 75G
Fine-tuning 66G 69G
Table 5: Memory usage of GPU during pre-training and fine-tuning stages.
GPU Memory Overhead of EE-MLLM
We assess the GPU memory overhead of EE-MLLM during both the pre-training and fine-tuning stages. Similarly, we compare EE-MLLM with LLaVA at a resolution of $336\times 336$. We perform experiments on 8$\times$×H800, with a global batch size of 256 for pre-training, 128 for fine-tuning. The comparison of memory usage is detailed in Tab 5. During the pre-training stage, EE-MLLM exhibits significant lower memory usage, at 32G, while the memory usage of LLaVA is 75G. During the fine-tuning stage, as the primary memory usage is focused on the trainable LLM, the memory usage advantage becomes less noticeable, with EE-MLLM consuming 66G and LLaVA-v1.6 utilizing 69G.
Ablation Study
Implementation Details
Following LLaVA-v1.5 (Liu et al. 2023a), we employ Vicuna-7b-v1.5 (Zheng et al. 2023) as our foundational LLM. The training data is consistent with LLaVA-v1.5, encompassing both pre-training data and supervised fine-tuning data.
Comparisons with LLaVA
In Tab. 4, we compare EE-MLLM with LLaVA under the same settings across both general and fine-grained benchmarks. For the resolution of $336\times 336$, we employ CLIP-ViT-L-14 as the vision encoder, ensuring complete alignment with LLaVA-v1.5. EE-MLLM demonstrates comparable performance to LLaVA on general benchmarks. In terms of fine-grained benchmarks, at $336\times 336$ resolution, EE-MLLM exhibits promising performance on AI2D and ChartQA but slightly underperforms LLaVA on OCRBench and TextVQA. The average score for the $336\times 336$ resolution is 47.1, which is 98% of LLaVA’s 48.1. For the resolution of $672\times 672$, we utilize SigLIP as the vision encoder. As shown in Tab. 4, EE-MLLM obtains comparable results on AI2D and TextVQA, and the average score maintains 98% of LLaVA’s performance.
V O FFN SeedB Hall TextVQA ChartQA DocVQA Avg.
$336\times 336$
65.3 24.4 44.8 14.4 18.4 33.5
66.3 25.9 45.5 16.7 19.2 34.7
65.4 25.3 45.4 15.6 19.1 34.2
65.8 25.1 43.5 14.8 17.9 33.4
66.1 25.7 46.1 15.8 18.6 34.5
$672\times 672$
66.1 25.2 56.8 19.7 28.6 39.3
66.6 25.5 58.2 23.0 29.8 40.6
66.1 24.3 58.2 21.6 30.1 40.1
66.4 24.1 56.3 18.5 26.5 38.4
66.9 27.6 58.8 24.0 31.8 41.8
Table 6: Ablation of the aligner. Checkmarks indicate the weights employed in the aligner.
Ablation of the Aligner
In Tab. 6, we evaluate the aligner variants with different mapping weights. We employ SigLIP initialized from Idefics2 (Laurençon et al. 2024) to conduct experiments at different input resolutions. We remove the different weights from the aligner and train the model at a resolution of $336\times 336$. Checkmarks indicate the weights employed in the aligner. We have three findings: 1) Removing the entire aligner, as depicted in the first row, significantly reduces performance across multiple benchmarks. Specifically, the TextVQA score drops from 46.1 to 44.8, and the average score for the five benchmarks decreases from 34.5 to 33.5. This result highlights the aligner’s effectiveness in aligning visual features with the LLM feature space, allowing text tokens to capture crucial visual information and address questions through causal cross-attention modules. 2) When ablating individual weights in the aligner, we find that maintaining the structure is more critical. The absence of $V$ or $O$ has a relatively minor impact on low-resolution input, even resulting in slightly higher average performance when $V$ is missing. However, when the FFN is missing, the aligner’s structure no longer resembles a transformer block, leading to a significant performance loss. 3) We directly increase the input image resolution to $672\times 672$ without additional training and compare the variants with different aligner types. We observe that, at high-resolution input, the absence of $V$ or $O$ weights results in a substantial decline in fine-grained benchmarks such as TextVQA, ChartQA, and DocVQA. This finding demonstrates the importance of a complete aligner when applied to high-resolution images.
Visualization
We present four examples sampled from BLINK (Fu et al. 2024) and RealWorldQA (XAI 2024) to assess the impact of the architectural changes in Fig 3. The first example demonstrates that EE-MLLM can perceive fine-grained visual context within the image, such as the color of traffic lights. The second and third examples highlight EE-MLLM’s ability to comprehend the position of objects. Specifically, EE-MLLM can accurately identify guitar’s position relative to the table and the dog’s position. The last example reveals that EE-MLLM can distinguish subtle differences in visual content.
Conclusion
In this paper, we revisit previous studies on multimodal large language models and categorize them into two groups: self-attention-based methods and cross-attention-based methods. The former is data-efficient but computationally inefficient, while the latter is compute-efficient but data-inefficient. To ensure data-efficient while maintaining compute-efficient, we propose the composite attention mechanism for EE-MLLM, which incorporates the composite attention module for computational efficiency and the aligner for data efficiency. We conduct comprehensive experiments on general benchmarks and fine-grained benchmarks and find that EE-MLLM achieves state-of-the-art performance on most benchmarks. We also assess the inference speed of EE-MLLM in real-world scenarios, and the results demonstrate that EE-MLLM holds a significant advantage in inference.

### v3

@@
 Abstract
-In the realm of multimodal research, numerous studies leverage substantial image-text pairs to conduct modal alignment learning, transforming Large Language Models (LLMs) into Multimodal LLMs and excelling in a variety of visual-language tasks.
-The prevailing methodologies primarily fall into two categories: self-attention-based and cross-attention-based methods.
-While self-attention-based methods offer superior data efficiency due to their simple MLP architecture, they often suffer from lower computational efficiency due to concatenating visual and textual tokens as input for LLM.
-Conversely, cross-attention-based methods, although less data-efficient due to additional learnable parameters, exhibit higher computational efficiency by avoiding long sequence input for LLM.
-To address these trade-offs, we introduce the Data-Efficient and Compute-Efficient Multimodal Large Language Model (EE-MLLM).
-Without introducing additional modules or learnable parameters, EE-MLLM achieves both data and compute efficiency.
+Recent advancements in Multimodal Large Language Models (MLLMs) have demonstrated satisfactory performance across various vision-language tasks.
+Current approaches for vision and language interaction fall into two categories: self-attention-based and cross-attention-based methods.
+However, both approaches present inherent limitations, forcing a trade-off between data and computational efficiency.
+To address this issue, we introduce the Data-Efficient and Compute-Efficient MLLM (EE-MLLM).
 Specifically, we modify the original self-attention mechanism in MLLM to a composite attention mechanism.
-This mechanism has two key characteristics: 1) Eliminating the computational overhead of self-attention within visual tokens to achieve compute efficiency, and 2) Reusing the weights on each layer of LLM to facilitate effective modality alignment between vision and language for data efficiency.
+This mechanism has two key characteristics: 1) eliminating the computational overhead of self-attention among visual tokens to achieve compute efficiency, and 2) reusing the weights from each layer of LLM to facilitate effective vision-language modality alignment for data efficiency.
+As a result, EE-MLLM significantly outperforms Flamingo with limited training data, and reduces the prefilling time to 79 ms on an H800 GPU, compared to LLaVA’s 277 ms.
+To further investigate the efficiency of EE-MLLM, we present a training-free variant named EE-MLLM-F, which reduces the computation cost of self-attention-based method without additional training.
 Experimental results demonstrate the effectiveness of EE-MLLM across a range of benchmarks, including general-purpose datasets like MMBench and SeedBench, as well as fine-grained tasks such as TextVQA and DocVQA.
@@
 (3) EE-MLLM introduces a composite attention mechanism to eliminate the computational overhead of self-attention within visual tokens and reuse the weights as aligners on each layer of LLM to facilitate modality alignment.
-Introduction
-Recently, Large Language Models (LLMs) have gained significant attention due to their surprising performance on various natural language tasks.
-However, real-world scenarios often involve more than just linguistic modality, making it crucial to extend LLM to Multimodal LLMs.
-The key to expansion is to perform modality alignment, that is, to learn to map the remaining modalities with the same semantics to their corresponding language modality in the feature space of the pre-trained LLM.
-In the field of multimodal LLM research involving vision and language, a series of research works on modality alignment have been carried out and achieved excellent performance.
-These methods can be roughly divided into two categories: self-attention-based methods and cross-attention-based methods.
-The self-attention-based methods firstly leverage a simple MLP layer to project the visual tokens into the input space of LLM, and then concatenate the visual tokens with the textual tokens, which is shown in Fig.1 (1).
-These flattened concatenated long token sequences are fed into the LLM to further model the dependencies of various tokens.
-These methods exhibit great data efficiency due to two factors: 1) The alignment module is simple and has a small number of learnable parameters.
-2) Visual tokens aligned in the input space of LLM will naturally lead to the alignment between vision and language in each layer of LLM.
-Therefore, the training procedure does not need to focus on the modality alignment within the LLM, enabling data efficiency.
-For example, LLaVA (Liu et al.
-2023b) achieves promising performance by pre-training on only 558k data and fine-tuning on only 665k instruction data.
-However, the flattening operation of directly concatenating visual tokens and textual tokens expands the length of the token sequence from $T$ to $V+T$ (where $T$ is the number of textual tokens, $V$ is the number of visual tokens), and the computational cost of the language model layer (self-attention structure) increases from $O(T^{2})$ to $O((V+T)^{2})$, results in lower computer-efficiency.
-Especially when the input image is high-resolution, the number of visual tokens increases, and the compute inefficiency will prolong the training and inference time.
-Specifically, LLaVA (Liu et al.
-2023b) and Deepseek-VL (Lu et al.
-2024) have 576 visual tokens, Sphinx (Lin et al.
-2023) utilizes 2,890 visual tokens.
-The cross-attention-based methods directly insert new cross-attention layers between existing pre-trained language model layers.
-As shown in Fig.1 (2), the keys and values in these layers are obtained from the visual tokens while the queries are derived from the text tokens.
-These methods have higher compute efficiency because the length of the input sequence is always kept as $T$ and does not increase with the increase of visual tokens.
-However, these methods usually require a large amount of pre-training data, leading to data inefficiency.
-For example, Flamingo (Alayrac et al.
-2022) is pre-trained on 1.8 billion image-text pairs (Jia et al.
-2021) and interleaved data.
-EVLM (Chen et al.
-2024) is pre-trained on 2.5 billion image-text pairs and 50 million interleaved image-text data.
-This data inefficiency arises from two factors: 1) The additional cross-attention layers will introduce a large number of learnable parameters, necessitating more training data.
-2) The training procedure has to focus on the modality alignment between vision and language on different layers of LLM, posing a huge optimization complexity for modality alignment.
-By reviewing existing methods, we conclude that the key points to ensure data efficiency and compute efficiency are 1) Avoiding an increase in the length of input sequence for LLM, 2) The modality alignment module should have fewer parameters, and the alignment needs to be effective.
+1 Introduction
+Multimodal Large Language Models (MLLMs) have gained significant attention due to their remarkable performance on diverse vision-language tasks, including image captioning, Visual Question Answering (VQA) and so on.
+Currently, mainstream MLLMs can be categorized into self-attention-based and cross-attention-based methods.
+However, these two predominant approaches entail a trade-off between compute and data efficiency.
+In general, cross-attention-based methods excel in computational efficiency but require substantial training data for effective modality alignment, thus exhibiting low data efficiency.
+Conversely, self-attention-based methods have high data efficiency, which can achieve modality alignment with limited data, but often suffer from high computational complexity, resulting in reduced computational efficiency.
+Specifically, self-attention-based methods project visual tokens into LLM’s input space with a simple MLP and concatenate them with text tokens, as illustrated in Fig.1 (1).
+Their data efficiency stems from: 1) Few learnable parameters in the alignment module, and 2) When visual tokens are aligned to the input space via a projector during input, these remain consistent with the processing space of each LLM’s decoder layer.
+Therefore, LLaVA achieves strong performance with only 558k pre-training and 665k instruction-tuning samples [28], in other words, it does not require a large amount of data for modality alignment.
+Cross-attention-based methods insert cross-attention layers between existing LLM layers.
+As shown in Fig.1 (2), the keys and values in these layers are derived from visual tokens, while the queries come from text tokens.
+These methods introduce a large number of learnable parameters, which require substantial training data, leading to data inefficiency.
+For example, Flamingo is pre-trained on 1.8 billion image-text pairs [1, 16] and interleaved data.
+Thus, this data inefficiency arises from two factors: 1) Learnable parameters from additional cross-attention layers, and 2) The training procedure has to focus on the modality alignment between vision and language on different layers of LLM, posing a huge optimization complexity for modality alignment.
+Regarding compute efficiency, self-attention methods concatenate visual and text tokens, which increases sequence length from $T$ to $V+T$, raising computational complexity of LLM from $O(T^{2})$ to $O((V+T)^{2})$.
+In contrast, cross-attention methods can achieve high compute efficiency because the input sequence length remains fixed at $T$ and does not increase with the number of visual tokens.
+By reviewing existing methods, we conclude that the key points to ensure compute efficiency and data efficiency are 1) Avoiding an increase in the length of input sequence for LLM, 2) The modality alignment module should have fewer parameters, and the alignment needs to be effective.
 Taking two summary viewpoints into account, we present a data-Efficient and compute-Efficient Multimodal Large Language Model (EE-MLLM).
-Specifically, we propose a composite attention mechanism to replace the original self-attention mechanism in MLLM.
+Specifically, we propose a composite attention mechanism to replace the original self-attention mechanism in LLM.
 The composite attention mechanism has two features: 1) Eliminating the computational overhead of self-attention within visual tokens to achieve compute efficiency.
-The original self-attention mechanism of MLLM consists of three components: self-attention within visual tokens, self-attention within text tokens, and cross-attention between text and visual tokens.
-We observe that the self-attention within visual tokens is redundant.
-This is because a) the interaction between visual tokens is already well-processed by the vision encoder, and b) LLM can facilitate the implicit interaction among visual tokens via the information aggregation property of LLM (Wang et al.
-2023b).
-Therefore, we eliminate the self-attention within visual tokens and maintain the length of the input sequence for LLM as the length of text tokens, thereby facilitating computational efficiency.
+Similar to FastV [9], we observe that the self-attention within visual tokens is redundant.
+This redundancy stems from: a) the interaction between visual tokens is already well-processed by the vision encoder, and b) LLM can facilitate the implicit interaction among visual tokens via the information aggregation property of LLM [42].
+Therefore, we eliminate self-attention among visual tokens and maintain the length of the input sequence for the LLM to be the same as that of text tokens, thereby facilitating computational efficiency.
 2) Instead of introducing additional modules like cross-attention-based methods, we reuse the weights on each layer of LLM to facilitate effective modality alignment between vision and language for data efficiency.
@@
 This means the training procedure does not need to focus on the modality alignment within the LLM, which enables data efficiency.
-To demonstrate the effectiveness of EE-MLLM, we evaluate its performance on both general benchmarks and fine-grained benchmarks, such as MMBench-EN, MME, ScienceQA, AI2D, TextVQA, and DocVQA.
-In comparison with state-of-the-art MLLMs, our EE-MLLM exhibits promising results on both general and fine-grained benchmarks.
-Simultaneously, EE-MLLM proves to be computationally efficient in the inference stage.
-When applied to high-resolution image inputs, EE-MLLM maintains comparable performance at a significantly reduced computational cost.
-For instance, with a $980\times 980$ input image, the number of visual tokens can reach 4,900, and the FLOPs of EE-MLLM are $70\%$ of those in self-attention-based methods.
-We also assess the inference speed of EE-MLLM in a real-world scenario, confirming its efficiency.
-Specifically, when the number of generated tokens is set to 32, EE-MLLM achieves an inference speed of 77 tokens per second on a single NVIDIA H800 GPU, which is 1.9 times faster than self-attention-based methods at the same resolution $980\times 980$.
+With the proposed composite attention, EE-MLLM can achieve data-efficient and compute-efficient at the same time.
+We also investigate a training-free variant of EE-MLLM where we directly apply the composite attention mechanism to select layers of LLaVA, thereby reducing computational cost without requiring additional training.
+Notably, this approach maintains comparable performance, demonstrating computational efficiency without sacrificing accuracy.
+We evaluate EE-MLLM on both general benchmarks and fine-grained benchmarks and our model demonstrates promising results across all these evaluations while maintaining high computational efficiency during inference.
+For a $980\times 980$ input image, EE-MLLM requires only $76\%$ of the FLOPs needed by self-attention methods.
+Moreover, in real-time scenarios that require high response speed but fewer output tokens, EE-MLLM can significantly reduce the prefilling time.
+On a single NVIDIA H800 GPU, EE-MLLM takes only 79 ms for prefilling with one input image, whereas LLaVA consumes 277 ms.
 Figure 2: Our composite attention mechanism consists of the composite attention module and the aligner.
@@
   1. 1.
-    We revisit modality alignment in multimodal LLMs, pointing out the issue of efficiency and efficacy in self-attention-based and cross-attention-based methods.
+    We propose composite attention mechanism for EE-MLLM, which enhances both data and compute efficiency by eliminating redundant attention calculation among visual tokens and reusing existing LLM weights for modality alignment.
   2. 2.
-    We propose a composite attention mechanism for EE-MLLM to enhance the data and computational efficiency.
+    We also introduce a training-free variant EE-MLLM-F, which adapts the composite attention mechanism to existing self-attention-based models, reducing computational costs without requiring additional training while preserving performance.
   3. 3.
-    Our EE-MLLM achieves outstanding performance on a variety of benchmarks, while the inference speed is much improved.
-Related Work
+    Experimental results demonstrate that EE-MLLM achieves superior performance across various benchmarks, while requiring only 76% of the FLOPs compared to self-attention methods and significantly reducing the prefilling time from 277 ms to 79 ms.
+2 Related Work
 Self-attention-based methods.
-Self-attention-based methods (Dai et al.
-2024; Zhu et al.
-2024; Chen et al.
-2023a; Bai et al.
-2023; Chen et al.
-2023b; Fang et al.
-2023) employ a projector to align visual features with the input space of LLM.
-LLaVA (Liu et al.
-2023b, a) uses a simple MLP layer to connect the vision encoder of CLIP (Radford et al.
-2021) with Vicuna (Zheng et al.
-2023).
-DeepSeek-VL (Lu et al.
-2024) employs hybrid vision encoders, including SigLIP (Zhai et al.
-2023) for high-level semantic features and SAM-B (Kirillov et al.
-2023) for low-level features.
-Considering the scenarios that require fine-grained information, such as document recognition and form understanding, recent work extends MLLM to support high-resolution inputs.
-LLaVA-Next (Liu et al.
-2024) proposes dividing the image into small patches for extracting local features and concatenating the downsampled image feature as global context.
-Monkey (Li et al.
-2024b) also partitions high-resolution images into small patches, processing each patch independently with a trainable resampler.
-Since the self-attention-based methods are compute-inefficiency, the overhead for high-resolution inputs can be more severe.
-Deco (Yao et al.
-2024) and HoneyBee (Cha et al.
-2024) carefully design projectors to reduce the number of visual tokens with minimal information loss.
+Self-attention-based methods [10, 48, 6, 3, 8, 12, 33] employ a projector to align visual features with the input space of LLM.
+LLaVA [28, 27] uses a simple MLP layer to connect the vision encoder of CLIP [39] with Vicuna [47].
+Considering the scenarios that require fine-grained information, such as document recognition and form understanding, recent work extends MLLM to support high-resolution inputs, resulting in long input sequences with numerous visual tokens.
+LLaVA-Next [29] proposes dividing the image into small patches for extracting local features and concatenating the downsampled image feature as global context.
+Monkey [26] also partitions high-resolution images into small patches, processing each patch independently with a trainable resampler.
+Another line of research focuses on compressing the number of visual tokens to reduce computational cost.
+Deco [44] and HoneyBee [5] carefully design projectors to reduce the number of visual tokens with minimal information loss.
 However, these methods still compromise fine-grained capabilities when reducing the number of visual tokens.
@@
 Cross-attention-based methods.
-Cross-attention-based methods (Laurençon et al.
-2023; Alayrac et al.
-2022; Awadalla et al.
-2023; Chen et al.
-2024) typically incorporate additional cross-attention modules into the decoder layers of LLM to integrate visual information without extending the length of the text token sequence.
-Flamingo (Alayrac et al.
-2022) introduces a novel gated xattn-dense block, incorporating a tanh-gating mechanism for training stability.
+Cross-attention-based methods [22, 1, 2, 7] typically incorporate additional cross-attention modules into the decoder layers of LLM to integrate visual information without extending the length of the text token sequence.
+Flamingo [1] introduces a novel gated xattn-dense block, incorporating a tanh-gating mechanism for training stability.
 Flamingo is trained on large-scale interleaved image-text data and supports in-context few-shot learning.
-EVLM (Chen et al.
-2024) enhances Flamingo by: (1) utilizing hierarchical ViT features without a resampler, (2) replacing the image token with a set of learnable tokens in the text input, and (3) introducing MoE to boost performance.
+EVLM [7] enhances Flamingo by: (1) utilizing hierarchical ViT features without a resampler, (2) replacing the image token with a set of learnable tokens in the text input, and (3) introducing MoE to boost performance.
 Although these cross-attention-based methods are computationally efficient, the additional cross-attention blocks lead to a large number of trainable parameters and a heavy reliance on pre-training data.
 Our EE-MLLM avoids introducing any extra trainable parameters and repurposes the weights of LLM to map visual tokens into the LLM feature space, thus achieving promising performance with a small amount of training data.
-Method
-Model Architecture
-EE-MLLM comprises a visual encoder, a two-layer MLP serving as the projector, and the LLM with composite decoder layers.
-We propose a composite attention mechanism and design a composite decoder layer for EE-MLLM to achieve data efficiency and compute efficiency.
-The composite decoder layer consists of an aligner and a composite attention module.
-Composite Attention Module.
-The original self-attention mechanism comprises self-attention within visual tokens, self-attention within text tokens, and cross-attention between text tokens and visual tokens.
-We observe that self-attention within visual tokens is redundant for two reasons: 1) The interaction between visual tokens is already well-learned in the vision encoder, and 2) LLM can enable the implicit interaction between visual tokens via the information aggregation property (Wang et al.
-2023b).
-Therefore, we eliminate self-attention within visual tokens and obtain the composite attention module.
+3 Method
+3.1 Model Architecture
+EE-MLLM is composed of three main components: a visual encoder, a projector made up of two-layer MLPs, and an LLM with composite decoder layers.
+The composite decoder layer contains an aligner and a composite attention module, designed to ensure data and compute efficiency.
+3.1.1 Composite Attention Module
+The original self-attention mechanism builds correlations in three aspects: intra-interaction among visual tokens, intra-interaction among text tokens, and inter-interaction among visual tokens and text tokens.
+In decoder layers, we observe that the process with visual tokens is redundant for two reasons: 1) The interaction among visual tokens is already well-learned in the vision encoder, and 2) the visual information has been aggregated into the text tokens through the inter-interaction among text tokens and visual tokens [42].
+Therefore, we eliminate intra-interaction among visual tokens and obtain the composite attention module.
 We assume that visual tokens $I\in\mathbb{R}^{k\times h}$ and text tokens $T\in\mathbb{R}^{n\times h}$, where $k$ and $n$ are the lengths of visual tokens and text tokens, respectively.
@@
 As illustrated in Fig.
-2, our composite attention module uses text tokens as the queries and the concatenation of visual tokens and text tokens as keys and values for cross-attention.
+2, our composite attention module uses text tokens as queries and the concatenation of visual tokens and text tokens as keys and values for cross-attention.
 Specifically, $X^{q}=TW_{q}$, $X^{k}=[I;T]W_{k}$, and $X^{v}=[I;T]W_{v}$.
@@
 The attention mask for self-attention is a lower triangular matrix, whereas the attention mask for the composite attention module is a trapezoidal matrix.
-This implies that text tokens will attend to all previous tokens, including visual tokens, while the interaction within visual tokens is eliminated.
+This implies that text tokens will attend to all previous tokens, including visual tokens, while intra-interaction among visual tokens is eliminated.
+3.1.2 Aligner
+We introduce the aligner, which leverages the existing weights at each layer of LLM to enhance modality alignment between vision and language without the need for additional modules such as cross-attention-based methods.
+By reusing these weights, visual tokens aligned in the input space of the LLM can naturally facilitate modality alignment at each layer.
+This approach eliminates the necessity for the training procedure to concentrate on modality alignment within the LLM, thereby promoting data efficiency.
+As depicted in Fig.
+2, in each layer, text tokens are mapped by the value matrix ($W_{v}$), output matrix ($W_{o}$), and FFN.
+Hence, we make visual tokens follow the same mapping relationship in the aligner.
+Specifically, the aligner comprises $W_{v}$, $W_{o}$, and FFN, which share weights with the corresponding LLM modules.
+The residual connection is also implemented in the aligner.
+The aligner can be formulated as follows:
+$O=\mathrm{FFN}(IW_{v}W_{o}+I)+IW_{v}W_{o}+I$ (2)
+This part of the decoder layer skips attention calculation.
+3.2 Training-Free Adaptation of EE-MLLM
+Since our composite attention mechanism does not introduce additional parameters for the interaction between vision and language, which reuses the weights of each LLM layer for modality alignment, we can directly apply this method to self-attention-based models without requiring additional training.
+Specifically, given a well-trained MLLM with $L$ layers, we can replace the self-attention blocks in selected layers $\mathcal{L}_{\text{replace}}\subseteq\{1,2,...,L\}$ with our composite attention.
+For each layer $l\in\mathcal{L}_{\text{replace}}$, the original self-attention is substituted with our composite attention, and the weights of $W_{v}$, $W_{o}$, and FFN are reused to align the visual tokens.
+Inspired by FastV [9], which demonstrates the attention of visual tokens becomes sparse in deep layers, we maintain the self-attention mechanism in the early layers of LLM while replacing the deeper layers with our composite attention.
+With this adaption, we can reduce the computation cost of the self-attention-based MLLM in deep layers, while maintaining strong performance.
+Model LLM MMB-T MME ScienceQA HallB MMMU CCBench SeedBench BLINK
+InstructBLIP Vicuna-7B 36.0 1137.1 54.7 31.2 30.6 12.7 44.5
+MiniGPT-4-v1 Vicuna-7B 12.2 770.6 39.0 31.9 23.6 1.8 31.6
+MiniGPT-4-v2 Llama2-7B 24.3 708.4 54.1 30.0 25.0 1.4 29.4
+Idefics-instruct Llama-7B 48.2 942.0 51.6 27.3 18.4 7.8 45.0 38.3
+OpenFlamingo v2 MPT-7B 6.6 535.0 45.7 29.4 28.2 6.3 28.8
+Qwen-VL Qwen-7B 38.2 334.1 57.7 29.9 29.6 6.1 52.5 27.9
+Qwen-VL-Chat Qwen-7B 60.6 1467.8 65.5 36.8 37.0 41.2 64.8 28.2
+VLoRA Vicuna-7B 63.4 1311.3 66.4 26.4 36.0 28.6
+ShareGPT4V Vicuna-7B 64.6 1561.4 68.2 28.6 37.2 30.8 69.3 40.9
+LLaVA-v1.5 Vicuna-7B 62.3 1510.7 66.8 27.6 35.7 27.5 65.8 39.7
+LLaVA-v1.6 Vicuna-7B 66.5 1475.6 68.5 27.6 37.6 24.3 69.6 41.6
+EE-MLLM Vicuna-7B 70.4 1528.1 77.7 38.6 33.4 37.3 70.2 43.2
+Table 1: Comparison with state-of-the-art methods on general benchmarks, including MMBench, MME, ScienceQA, HallusionBench, MMMU, CCBench, SeedBench, and BLINK.
+MMB-T indicates we report the results on MMBench-TEST-EN-V11.
+Model LLM AI2D OCRBench TextVQA ChartQA Seed2 Plus
+InstructBLIP Vicuna-7B 40.6 27.6 33.6 10.9 29.5
+MiniGPT-4-v1 Vicuna-7B 28.4 17.2 15.2
+MiniGPT-4-v2 Llama2-7B 30.5 3.1 23.3
+Idefics-instruct Llama-7B 42.2 25.2 35.4
+OpenFlamingo v2 MPT-7B 31.7 14.9 16.3 28.7
+Qwen-VL Qwen-7B 57.7 12.7 63.1 59.0 40.1
+Qwen-VL-Chat Qwen-7B 63.0 48.8 60.7 49.8 40.6
+ShareGPT4V Vicuna-7B 58.0 37.1 51.1 21.3 46.1
+LLaVA-v1.5 Vicuna-7B 55.5 31.8 45.5 17.8 41.3
+LLaVA-v1.6 Vicuna-7B 67.0 53.2 64.4 55.4 51.6
+EE-MLLM Vicuna-7B 65.7 57.4 69.0 68.4 53.5
+Table 2: Comparison with state-of-the-art methods on fine-grained benchmarks, including AI2D, OCRBench, TextVQA, ChartQA, and SeedBench-2 Plus.
+3.3 Analysis of the Computational Overhead
+EE-MLLM reduces computational overhead without introducing extra parameters.
+This section analyzes the compute efficiency of EE-MLLM compared to LLaVA in terms of FLOPs.
+We consider an LLM with $d$ blocks, hidden state dimension $h$, intermediate size $m$, $T$ text tokens, and $V$ visual tokens.
+The total FLOPs of LLaVA are given by:
+$\text{FLOPs}_{\text{LLaVA}}=\underbrace{8(V+T)dh^{2}+4(V+T)^{2}dh}_{\text{self-attention module}}+\underbrace{6(V+T)dhm}_{\text{FFN module}}.$ (3)
+The total FLOPs of EE-MLLM are given by:
+$\begin{split}\text{FLOPs}_{\text{EE-MLLM}}&=\underbrace{(6V+8T)dh^{2}+(2T^{2}+2VT)dh}_{\text{composite attention module}}\\ &\quad+\underbrace{6(V+T)dhm}_{\text{FFN module}}.\end{split}$ (4)
+For a typical training scenario with a $980\times 980$ input image (yielding 4,900 visual tokens) and 256 text tokens, EE-MLLM reduces computational requirements by approximately 24% compared to LLaVA, demonstrating significant efficiency gains.
+4 Experiments
+4.1 Implementation Details
+Model Configurations.
+We employ Vicuna-7b-v1.5 [47] as our LLM and SigLIP [46] as the vision encoder.
+Specifically, SigLIP is initialized from Idefics2 [21], which supports dynamic resolutions up to 980 $\times$ 980.
+The projector consists of a two-layer MLP.
+Pre-training Configurations.
+During the pre-training stage, we follow LLaVA to adopt blip-558k [25] for training one epoch.
+We freeze the LLM and the vision encoder, making only the projector trainable.
+We adopt AdamW [32] optimizer.
+The learning rate is 1e-3, followed by a linear warm-up scheduler and then a cosine decay scheduler.
+The global batch size is 256.
+Fine-tuning Configurations.
+During the fine-tuning stage, we freeze the vision encoder and update the weights of the LLM and projector.
+The total amount of supervised fine-tuning data is 3 million samples.
+The sources of fine-tuning data are provided in Appendix 8.1.
+We utilize the AdamW optimizer with a global batch size of 128 and a learning rate of 2e-5.
+The learning rate scheduler is the same as the pre-training stage.
+Training is conducted on 16 H800 GPUs.
+4.2 Evalution Benchmarks
+We conduct our evaluations using the VLMEvalKit [11], and the results of other state-of-the-art models are obtained from the same source.
+General Benchmarks.
+For general benchmarks, our evaluation includes MMBench-EN [30], MME [13], ScienceQA [34], HallusionBench [15], MMMU [45], CCBench [30], SeedBench [23], and BLINK [14].
+Among them, MMBench-EN is a comprehensive multimodal benchmark specifically designed to assess the performance of multimodal large language models (MLLMs).
+It contains more than 3,000 multiple-choice questions spanning 20 distinct ability categories.
+We evaluate EE-MLLM on MMBench-EN-V1.1.
+Further details on the other general benchmarks are provided in Appendix 9.1.
+Fine-grained Benchmarks.
+We also compare EE-MLLM with other models on AI2D [19], OCRBench [31], TextVQA [40], ChartQA [36], DocVQA [37], and SeedBench 2 Plus [24].
+More details of these fine-grained benchmarks are provided in Appendix 9.2.
+Figure 3: Comparison of prefilling time between EE-MLLM and LLaVA.
+The X-axis represents the number of input images, where each image has a resolution of $980\times 980$.
+The Y-axis indicates the prefilling time.
+4.3 Comparisons with State-of-the-art Models
+General Benchmarks.
+In Tab.
+1, we compare EE-MLLM with various MLLMs equipped similar LLM [29, 3, 35] on general benchmarks.
+These benchmarks evaluate the comprehensive ability of MLLMs, including understanding and perception.
+EE-MLLM achieves comparable performance with state-of-the-art MLLMs in general benchmarks.
+Specifically, EE-MLLM achieves a score of 70.4 in MMBench and 1528.1 in MME, scores that are notably higer compared to LLaVA-v1.6, which also support high-resolution images input.
+This suggests that EE-MLLM possesses comprehensive perception and reasoning capabilities.
+What’s more, EE-MLLM also achieves promising performance on CCBench and SeedBench.
+Fine-grained Benchmarks.
+In Tab.
+2, we evaluate EE-MLLM on fine-grained benchmarks.
+These benchmarks demand strong visual perception capabilities from MLLMs, as they require extracting fine-grained information from images to accurately answer questions.
+On TextVQA and ChartQA, EE-MLLM achieves impressive performance, 4.6 higher on TextVQA and 13.0 higher on ChartQA compared with LLaVA-v1.6, which also supports high resolution.
+In the OCRBench, specifically designed to assess the OCR capabilities of MLLM, EE-MLLM outperforms by 4.2 compared to LLaVA-v1.6.
+These results demonstrate that despite EE-MLLM significantly reducing the computational overhead associated with visual tokens, it effectively maintains the model’s fine-grained capabilities.
+Model MMB-VAL SeedBench ScienceQA HallB
+LLaVA-v1.5 67.3 65.1 67.0 25.7
+EE-MLLM-F 67.4 65.1 67.0 32.4
+AI2D OCRBench TextVQA ChartQA
+LLaVA-v1.5 56.4 31.6 46.1 17.9
+EE-MLLM-F 56.4 30.7 45.5 17.4
+Table 3: Performance comparison of LLaVA-v1.5 with our composite attention mechanism across general and fine-grained benchmarks.
+4.4 Comparison of Prefilling Time
+Since we eliminate the intra-interactions among visual tokens, EE-MLLM can significantly reduce the prefilling time, thus accelerating the generation of the first output token.
+This efficiency is crucial in real-time scenarios, particularly for tasks that only require a few output tokens.
+For example, in content moderation systems, MLLMs must rapidly classify social medias as safe or unsafe quickly to prevent harmful content exposure while maintaining platform throughput.
+Similarly, industrial quality control systems require MLLMs to swiftly provide pass/fail decisions.
+We conduct comparisons of prefilling time on a single NVIDIA H800.
+The resolution of the input image is set to $980\times 980$.
+We illustrate the time consumed for prefilling by EE-MLLM and LLaVA in Fig.
+3.
+With one input image, EE-MLLM costs 79 ms for prefilling, while LLaVA costs 277 ms.
+When the input is extended to two images, the time consumed by LLaVA severely increases to 1086 ms, but the prefilling time of EE-MLLM only rises to 158 ms.
+We also test the scenario where the number of input images is 3, and the gap between EE-MLLM and LLaVA becomes even larger.
+To further analyze the efficiency of EE-MLLM, we compare the overall inference speed and provide a detailed analysis in Appendix 10.
+4.5 Training-free Adaption of EE-MLLM
+In Tab 3, we present the results of directly applying our composite attention mechanism to LLaVA-v1.5-7b without additional training.
+Specifically, we replace the attention mechanism in the last 16 layers of LLaVA-v1.5-7b with our composite attention approach.
+And EE-MLLM-F performs comparably to the original LLaVA-v1.5 on MMBench, SeedBench, and ScienceQA, and even achieves better results on HallusionBench.
+For fine-grained benchmarks, we observe only a slight performance drop on OCRBench and TextVQA.
+These results demonstrate that our composite attention mechanism can directly adapt to self-attention-based methods without the need for training, while maintaining performance without significant degradation.
+Model MMB-VAL SeedBench ScienceQA HallB AI2D OCRBench TextVQA ChartQA Avg. Ratio
+Resolution: $336\times 336$
+LLaVA-v1.5 67.3 65.1 67.0 25.7 56.4 31.6 46.1 17.9 47.1 100%
+Flamingo 60.7 53.0 66.9 28.2 52.2 13.0 27.4 12.6 39.3 81.7%
+EE-MLLM 66.8 64.1 66.7 25.2 56.2 29.1 43.3 17.3 46.1 97.9%
+Resolution: $672\times 672$
+LLaVA-v1.5 70.1 68.0 69.4 27.2 58.3 39.8 61.7 27.9 52.8 100%
+Flamingo 64.2 55.6 69.4 28.1 53.9 16.5 36.7 19.8 43.0 81.4%
+EE-MLLM 68.2 67.3 69.1 25.9 57.5 37.0 60.1 25.1 51.3 97.2%
+Table 4: Comparisons with LLaVA and Flamingo under the same settings, including LLM, vision encoder and training data.
+5 Ablation Study
+5.1 Implementation Details
+Following LLaVA-v1.5 [27], we employ Vicuna-7b-v1.5 [47] as our LLM, and CLIP-ViT-L-14 as vision encoder.
+The training data is consistent with LLaVA-v1.5.
+To support high-resolution image input, we replace CLIP-ViT-L-14 with SigLIP [46], which is derived from Idefics2 [21] and supports high-resolution image input.
+To compare with cross-attention based methods, we reproduce Flamingo under the similar settings, the implementation details are provided in Appendix 8.2.
+5.2 Comparisons with LLaVA and Flamingo
+In Tab.
+4, we reproduce LLaVA and Flamingo under the same settings with EE-MLLM.
+We follow LLaVA to employ blip-558k for pre-training and adopt LLaVA-665k for fine-tuning.
+For the resolution of $336\times 336$, we employ CLIP-ViT-L-14 as the vision encoder, ensuring complete alignment with LLaVA-v1.5.
+EE-MLLM demonstrates comparable performance to LLaVA on general benchmarks.
+In terms of fine-grained benchmarks, at $336\times 336$ resolution, EE-MLLM exhibits promising performance on AI2D and ChartQA but slightly underperforms LLaVA on OCRBench and TextVQA.
+The average score for the $336\times 336$ resolution is 46.1, which is 97.9% of LLaVA’s 47.1.
+And the performance of Flamingo is significantly worse than that of EE-MLLM and LLaVA-v1.5, with an average score of only 39.3, which is 81.7% of LLaVA’s score, this underperformance can be attributed to Flamingo’s architecture.
+Flamingo introduces additional learnable parameters and requires alignment of visual tokens across different LLM layers, necessitating a large volume of training data, for which the BLIP-558K and LLaVA-665K datasets are insufficient.
+For the resolution of $672\times 672$, we utilize SigLIP as the vision encoder.
+As shown in Tab.
+4, EE-MLLM obtains comparable results on AI2D and TextVQA, and the average score maintains 97.2% of LLaVA’s performance.
+Similarly, Flamingo only achieves an average score of 43.0, which can be attributed to inadequate training data.
+Aligner Components Evaluation Benchmarks
+V O FFN SeedB Hall TextVQA ChartQA DocVQA Avg.
+Resolution: $336\times 336$
+65.3 24.4 44.8 14.4 18.4 33.5
+66.3 25.9 45.5 16.7 19.2 34.7
+65.4 25.3 45.4 15.6 19.1 34.2
+65.8 25.1 43.5 14.8 17.9 33.4
+66.1 25.7 46.1 15.8 18.6 34.5
+Resolution: $672\times 672$
+66.1 25.2 56.8 19.7 28.6 39.3
+66.6 25.5 58.2 23.0 29.8 40.6
+66.1 24.3 58.2 21.6 30.1 40.1
+66.4 24.1 56.3 18.5 26.5 38.4
+66.9 27.6 58.8 24.0 31.8 41.8
+Table 5: Ablation study of the aligner components V, O, and FFN.
+Checkmarks indicate employed weights.
+To reduce ablation costs, we adopt SigLIP with dynamic resolution from Idefics2, enabling direct evaluation at $672\times 672$ from models trained at $336\times 336$.
+Figure 4: The visualization results of examples sampled from BLINK and RealWorldQA. Answers are generated by our EE-MLLM.
+5.3 Ablation Study of the Aligner
+In Tab.
+5, we remove the different weights from the aligner and train the model at a resolution of $336\times 336$.
+Notably, to reduce the cost of ablation experiments, we adopt SigLIP with dynamic resolution support, derived from Idefics2, as our vision encoder.
+Thus, models trained at $336\times 336$ can be directly evaluated at a resolution of $672\times 672$.
+We have three findings:
+1) Removing the entire aligner, as depicted in the first row, significantly reduces performance across multiple benchmarks.
+Specifically, the TextVQA score drops from 46.1 to 44.8, and the average score decreases from 34.5 to 33.5.
+This result highlights the aligner’s effectiveness in aligning visual features with the LLM feature space, allowing text tokens to capture crucial visual information and address questions through causal cross-attention modules.
+2) When ablating individual weights in the aligner, we find that maintaining the structure is more critical.
+The absence of $V$ or $O$ has a relatively minor impact on low-resolution input, even resulting in slightly higher average performance when $V$ is missing.
+However, when the FFN is missing, the aligner’s structure no longer resembles a transformer block, leading to a significant performance loss.
+3) The importance of a complete aligner becomes even more pronounced at higher resolution ($672\times 672$).
+When $V$ or $O$ weights are removed, we observe substantial performance drops on detail-oriented benchmarks like TextVQA, with scores decreasing from 58.8 to 58.2, ChartQA from 24.0 to 21.6, and DocVQA reducing from 31.8 to 30.1.
+This demonstrates that all components are essential for effectively processing high-resolution visual inputs, where fine-grained details significantly impact performance.
+5.4 Ablation Study of Training-free Adaption
+In Tab.
+6, we explore different layer configurations for training-free adaptation of EE-MLLM.
+Using LLaVA-v1.5-7b with 32 layers as our baseline, we investigate various configurations of attention mechanism modification.
+The first configuration represents the original LLaVA, where all layers from 0 to 31 use the standard self-attention mechanism, achieving an average score of 35.5.
+When we apply the composite attention mechanism to layers 8 to 31, we observe a slight performance decrease to 32.4 on average, with improvements only on HallusionBench.
+Notably, the third configuration, which employs composite attention for layers 16 through 31 while maintaining self-attention for layers 0 through 15, achieves the best overall performance with a 36.5 average.
+This configuration not only maintains comparable performance on SeedBench, TextVQA, ChartQA, and DocVQA compared to the baseline, but also demonstrates significant improvement on HallucinationBench with a score of 32.4 versus 25.7.
+To further validate the effectiveness of our composite attention mechanism, we implement a configuration in the fourth row that maintains self-attention only in layers 0 through 15 and drops visual tokens in deeper layers.
+This results in a substantial performance degradation across fine-grained benchmarks including TextVQA, ChartQA, and DocVQA, with the average score dropping to 27.2.
+This finding confirms the composite attention mechanism plays a crucial role in preserving performance on tasks requiring detailed visual understanding, while simply dropping visual information in deeper layers is detrimental.
+These experimental findings clearly suggest that applying composite attention selectively to deeper layers from 16 to 31 provides an optimal balance between efficiency and overall performance across diverse multimodal tasks.
+Layer Configuration Evaluation Benchmarks
+Self-attn Layers Comp-attn Layers SeedB Hall TextVQA ChartQA DocVQA Avg.
+$[0,31]$ 65.1 25.7 46.1 17.9 22.5 35.5
+$[0,7]$ $[8,31]$ 62.8 29.0 39.8 16.1 14.5 32.4
+$[0,15]$ $[16,31]$ 65.1 32.4 45.5 17.4 21.9 36.5
+$[0,15]$ 65.1 29.2 17.6 13.8 10.5 27.2
+Table 6: Ablation study on layer configuration for training-free adaptation.
+“Self-attn Layers” indicates which layers maintain the original self-attention mechanism, while “Comp-attn Layers” shows which layers are modified to use the composite attention mechanism.
+The notation $[i,j]$ indicates layers from index $i$ to $j$.
+5.5 Visualization
+We present four examples sampled from BLINK [14] and RealWorldQA [43] to assess the impact of the architectural changes in Fig 4.
+The first example demonstrates that EE-MLLM can perceive fine-grained visual context within the image, such as the color of traffic lights.
+The second and third examples highlight EE-MLLM’s ability to comprehend the position of objects.
+Specifically, EE-MLLM can accurately identify guitar’s position.
+6 Conclusion
+In this paper, we introduce EE-MLLM, a novel framework that unifies data efficiency and compute efficiency through two key innovations: (1) a composite attention mechanism that eliminates redundant self-attention computations among visual tokens, reducing FLOPs by 24% compared to self-attention methods, and (2) a parameter-free aligner that reuses existing LLM weights to align visual and text modalities, minimizing training data requirements.
+EE-MLLM achieves state-of-the-art performance across various benchmarks.
+Crucially, it reduces inference latency from 277 ms to 79 ms, enabling real-time applications such as content moderation and quality control.
+The training-free variant (EE-MLLM-F) further demonstrates adaptability, maintaining performance while cutting computational costs by selectively applying composite attention to deeper layers.
+Supplementary Material
+Figure 5: Comparison of EE-MLLM and LLaVA-v1.6 response results on challenging samples.
+7 Visualization
+Fig.
+5 provides two examples comparing EE-MLLM with LLaVA-v1.6 [29].
+The first image, a handwritten coffee menu, prompts a question about the image’s fine-grained details.
+Our EE-MLLM successfully identifies the price of the Flat White listed in the image, providing the correct answer, whereas LLaVA-v1.6 does not.
+The second example requires the model to convert the image’s table into markdown format.
+Our observations indicate that EE-MLLM’s response is more accurate, while LLaVA makes mistakes in the columns labeled “Open Source” and “Data Source”.
+8 Implementation Details
+8.1 Supervised Fine-tuning Data
+The details of our supervised fine-tuning data are shown in Tab.
+7 Following Deepseek-VL [33], our supervised fine-tuning data incorporates open-source gpt4v datasets, including ShareGPT4V [8], LVIS-Instruct4V [41], LAION-GPT4V [20], and TextOCR-GPT4V [4].
+Moreover, to enhance the fine-grained capabilities of MLLM, we introduce DVQA [18], PlotQA [38], CLEVR [17], DocVQA [37], ChartQA [36], and ScienceQA [34] for fine-tuning.
+To preserve language abilities during fine-tuning, we utilize a large volume of text-only data from various sources.
+The text-only data encompasses various dataset types, such as mathematical reasoning, multi-round dialogues, logic comprehension, and code.
+Meanwhile, the in-house data primarily consists of OCR-related, table and diagram comprehension, and mathematical reasoning data.
+8.2 Implementation of Flamingo
+In the ablation study, we reproduced Flamingo.
+Like LLaVA-v1.5-7B, we used CLIP-ViT-L-14 as the vision encoder and Vicuna-7B as the LLM.
+We introduced cross-attention blocks every 4 layers into the LLM for inter-modal interaction.
+The cross-attention block includes attention operations and MLP.
+Specifically, we inserted cross-attention blocks at layers 3, 7, 11, 15, 19, 23, 27, and 31.
+The training setup was consistent with LLaVA, including training data, learning rate, and optimizer.
+In the ablation, we ensured a fair comparison.
 Type Dataset Ratio
@@
 Table 7: Details of supervised fine-tuning data.
-Aligner.
-We introduce the aligner, which leverages the existing weights at each layer of LLM to enhance modality alignment between vision and language without the need for additional modules such as cross-attention-based methods.
-By reusing these weights, visual tokens aligned in the input space of the LLM can naturally facilitate modality alignment at each layer.
-This approach eliminates the necessity for the training procedure to concentrate on modality alignment within the LLM, thereby promoting data efficiency.
-As depicted in Fig.
-2, in each layer, text tokens are mapped by the value matrix ($W_{v}$), output matrix ($W_{o}$), and FFN.
-Hence, we make visual tokens follow the same mapping relationship in the aligner.
-Specifically, the aligner comprises $W_{v}$, $W_{o}$, and FFN, which share weights with the corresponding LLM modules.
-The residual connection is also implemented in the aligner.
-The aligner can be formulated as follows:
-$H$ $=IW_{v}W_{o}+I$ (2)
-$O$ $={\rm FFN}(H)+H.$ (3)
-This can be viewed as a decoder block without attention.
-Analysis of the Computational Overhead
-Our EE-MLLM can reduce the computational overhead without introducing additional parameters.
-In this section, we analyze the compute efficiency of EE-MLLM from the perspective of FLOPs, comparing it with LLaVA.
-The LLM consists of $d$ blocks, with a hidden states dimension of $h$, an input text length of $T$, and $V$ visual tokens.
-Considering an LLM, we assume the input sequence length is $L$, the FLOPs of the self-attention module and FFN are $8Lh^{2}+4L^{2}h$ and $16Lh^{2}$.
-For self-attention-based methods, such as LLaVA, the length of the input sequence is $L=V+T$.
-Therefore, the FLOPs of LLaVA are $24(T+V)dh^{2}+4(T+V)^{2}dh$.
-For EE-MLLM, the FLOPs of composite attention modules are $(6T+2V)h^{2}+4VTh+4T^{2}h$.
-Given that we introduce aligners, the FLOPs of FFN are the same as those of LLaVA, which are $16(V+T)h^{2}$.
-The aligners introduce an additional matrix multiplication by $O$, adding $2Vh^{2}$ FLOPs.
-Therefore, the total FLOPs of EE-MLLM are $2(11T+10V)dh^{2}+4VTdh+4T^{2}dh$.
-Particularly, when the resolution of the input image is $980\times 980$, the number of visual tokens is 4,900, and the length of text tokens is 256, which is a typical length during training, the FLOPs of EE-MLLM amount to 70% of those of LLaVA.
-Experiments
-Implementation Details
-Model Configurations.
-We employ Vicuna-7b-v1.5 (Zheng et al.
-2023) as our LLM and SigLIP (Zhai et al.
-2023) as the vision encoder.
-Specifically, SigLIP is initialized from Idefics2 (Laurençon et al.
-2024), which supports dynamic resolutions up to 980 $\times$ 980.
-The projector consists of a two-layer MLP, identical to that of LLaVA (Liu et al.
-2023b).
-Model LLM MMB-T MME ScienceQA HallB MMMU CCBench SeedBench BLINK
-InstructBLIP Vicuna-7B 36.0 1137.1 54.7 31.2 30.6 12.7 44.5 -
-MiniGPT-4-v1 Vicuna-7B 12.2 770.6 39.0 31.9 23.6 1.8 31.6 -
-MiniGPT-4-v2 Llama2-7B 24.3 708.4 54.1 30.0 25.0 1.4 29.4 -
-Idefics-instruct Llama-7B 48.2 942 51.6 27.3 18.4 7.8 45.0 38.3
-OpenFlamingo v2 MPT-7B 6.6 535 45.7 29.4 28.2 6.3 28.8 -
-Qwen-VL Qwen-7B 38.2 334.1 57.7 29.9 29.6 6.1 52.5 27.9
-Qwen-VL-Chat Qwen-7B 60.6 1467.8 65.5 36.8 37.0 41.2 64.8 28.2
-VLoRA Vicuna-7B 63.4 1311.3 66.4 26.4 36.0 28.6 - -
-ShareGPT4V Vicuna-7B 64.6 1561.4 68.2 28.6 37.2 30.8 69.3 40.9
-LLaVA-v1.5 Vicuna-7B 62.3 1510.7 66.8 27.6 35.7 27.5 65.8 39.7
-LLaVA-v1.6 Vicuna-7B 66.5 1475.6 68.5 27.6 37.6 24.3 69.6 41.6
-EE-MLLM Vicuna-7B 70.4 1528.1 77.7 38.6 33.4 37.3 70.2 43.2
-Table 2: Comparsion with state-of-the-art methods on general benchmarks, including MMBench, MME, ScienceQA, HallusionBench, MMMU, CCBench, SeedBench, and BLINK.
-MMB-T indicates we report the results on MMBench-TEST-EN-V11.
-Pre-training Configurations.
-During the pre-training stage, we follow LLaVA to adopt blip-558k (Li et al.
-2022) for training one epoch.
-We freeze the LLM and the vision encoder, making only the projector trainable.
-We adopt AdamW (Loshchilov and Hutter 2018) optimizer.
-The learning rate is 1e-3, followed by a linear warm-up scheduler and then a cosine decay scheduler.
-The global batch size is 256.
-Fine-tuning Configurations.
-During the fine-tuning stage, we freeze the vision encoder and update the weights of LLM and projector.
-The total quantity of supervised fine-tuning data amounts to 3 million.
-As displayed in Tab.
-1, following Deepseek-VL (Lu et al.
-2024), our supervised fine-tuning data incorporates open-source gpt4v datasets, including ShareGPT4V (Chen et al.
-2023b), LVIS-Instruct4V (Wang et al.
-2023a), LAION-GPT4V (LAION 2023), and TextOCR-GPT4V (Carter 2024).
-Moreover, to enhance the fine-grained capabilities of MLLM, we introduce DVQA (Kafle et al.
-2018), PlotQA (Methani et al.
-2020), CLEVR (Johnson et al.
-2017), DocVQA (Mathew, Karatzas, and Jawahar 2021), ChartQA (Masry et al.
-2022), and ScienceQA (Lu et al.
-2022) for fine-tuning.
-To preserve language abilities during fine-tuning, we utilize a large volume of text-only data from various sources.
-The text-only data encompasses various dataset types, such as mathematical reasoning, multi-round dialogues, logic comprehension, and code.
-Meanwhile, the in-house data primarily consists of OCR-related, table and diagram comprehension, and mathematical reasoning data.
-The optimizer is AdamW with a global batch size of 128.
-The learning rate is 2e-5 and the learning rate scheduler is the same as the pre-training stage.
-The training is conducted on 16 NVIDIA H800 GPUs.
-Evalution Benchmarks
-We conduct our evaluations using the VLMEvalKit (Duan et al. 2024), and the results of other state-of-the-art models are obtained from the same source.
-General Benchmarks.
-1) MMBench-EN (Liu et al.
-2023c) is a comprehensive multimodal benchmark specifically designed to assess the performance of MLLMs.
+9 Evalution Benchmarks
+We conduct our evaluations using the VLMEvalKit [11], and the results of other state-of-the-art models are obtained from the same source.
+9.1 General Benchmarks
+1) MMBench-EN [30] is a comprehensive multimodal benchmark specifically designed to assess the performance of MLLMs.
 It comprises over 3,000 multiple-choice questions spanning 20 ability categories.
 We evaluate EE-MLLM on the MMBench-EN-V1.1.
-2) MME (Fu et al.
-2023) evaluates advanced MLLMs in terms of perception and cognition, encompassing a total of 14 subtasks.
-To minimize the impact of prompt engineering on MLLMs, MME’s instructions are designed to elicit simple binary responses, such as ”please answer yes or no.”.
+2) MME [13] evaluates advanced MLLMs in terms of perception and cognition, encompassing a total of 14 subtasks.
+To minimize the impact of prompt engineering on MLLMs, MME’s instructions are designed to elicit simple binary responses, such as “please answer yes or no.”.
 We report the results of the perception split of MME.
-3) ScienceQA (Lu et al.
-2022)is derived from elementary and high school science curricula.
+3) ScienceQA [34]is derived from elementary and high school science curricula.
 The questions in ScienceQA cover three subjects: natural science, language science, and social science.
-4) HallusionBench (Guan et al.
-2023) is designed for evaluating image-context reasoning, comprising 346 images paired with 1,129 questions crafted by human experts.
+4) HallusionBench [15] is designed for evaluating image-context reasoning, comprising 346 images paired with 1,129 questions crafted by human experts.
 HallusionBench takes into account both language hallucinations and visual illusions across a diverse range of topics.
-5) MMMU (Yue et al.
-2023) collects 11.5K multimodal questions from college exams, quizzes, and textbooks, covering six core disciplines, spanning 30 subjects and 183 subfields, including 30 heterogeneous image types.
-6) CCBench (Liu et al.
-2023c), developed by the MMBench team, is specifically designed to evaluate MLLMs in the domain of Chinese Culture.
-7) SeedBench (Li et al.
-2023) encompasses 19K multiple choice questions, covering 12 evaluation dimensions, including both image and video.
+5) MMMU [45] collects 11.5K multimodal questions from college exams, quizzes, and textbooks, covering six core disciplines, spanning 30 subjects and 183 subfields, including 30 heterogeneous image types.
+6) CCBench [30], developed by the MMBench team, is specifically designed to evaluate MLLMs in the domain of Chinese Culture.
+7) SeedBench [23] encompasses 19K multiple choice questions, covering 12 evaluation dimensions, including both image and video.
 We only use questions with images for evaluation.
-8) BLINK (Fu et al.
-2024) contains 14 visual perception tasks which pose significant challenges for current multimodal LLMs.
-Model LLM AI2D OCRBench TextVQA ChartQA DocVQA Seed2 Plus
-InstructBLIP Vicuna-7B 40.6 27.6 33.6 10.9 - 29.5
-MiniGPT-4-v1 Vicuna-7B 28.4 17.2 - - - 15.2
-MiniGPT-4-v2 Llama2-7B 30.5 3.1 - - - 23.3
-Idefics-instruct Llama-7B 42.2 25.2 - - - 35.4
-OpenFlamingo v2 MPT-7B 31.7 14.9 16.3 - - 28.7
-Qwen-VL Qwen-7B 57.7 12.7 63.1 59.0 - 40.1
-Qwen-VL-Chat Qwen-7B 63.0 48.8 60.7 49.8 - 40.6
-ShareGPT4V Vicuna-7B 58.0 37.1 51.1 21.3 - 46.1
-LLaVA-v1.5 Vicuna-7B 55.5 31.8 45.5 17.8 - 41.3
-LLaVA-v1.6 Vicuna-7B 67.0 53.2 64.4 55.4 - 51.6
-EE-MLLM Vicuna-7B 65.7 57.4 69.0 68.4 67.4 53.5
-Table 3: Comparsion with state-of-the-art methods on fine-grained benchmarks, including AI2D, OCRBench, TextVQA, ChartQA, DocVQA, InfoVQA, and SeedBench-2 Plus.
-Model MMB-VAL SeedBench ScienceQA BLINK AI2D OCRBench TextVQA ChartQA Avg.
-$336\times 336$
-LLaVA-v1.5 67.3 65.1 67.0 40.7 56.4 24.4 46.1 17.9 48.1
-EE-MLLM 66.8 64.1 66.7 38.4 56.2 23.6 43.3 17.3 47.1
-$672\times 672$
-LLaVA-v1.5 70.1 68.0 69.4 40.1 58.3 39.8 61.7 27.9 54.4
-EE-MLLM 68.2 67.3 69.1 40.3 57.5 37.0 60.1 25.1 53.1
-Table 4: Comparisons with LLaVA under the same settings, including LLM, vision encoder and training data.
-Fine-grained Benchmarks.
-1) AI2D (Kembhavi et al.
-2016) emphasizes diagram interpretation and reasoning, comprising 5,000 diagrams and 15,000 questions and answers.
-2) OCRBench (Liu et al.
-2023d) aims to facilitate the assessment of MLLM OCR capabilities, including 29 datasets.
-3) TextVQA (Singh et al.
-2019) consists of 45,336 questions and 28,408 images that necessitate reasoning about text for answering.
+8) BLINK [14] contains 14 visual perception tasks which pose significant challenges for current multimodal LLMs.
+Figure 6: Inference speed ratios of EE-MLLM and LLaVA for single and double image inputs.
+9.2 Fine-grained Benchmarks
+1) AI2D [19] emphasizes diagram interpretation and reasoning, comprising 5,000 diagrams and 15,000 questions and answers.
+2) OCRBench [31] aims to facilitate the assessment of MLLM OCR capabilities, including 29 datasets.
+3) TextVQA [40] consists of 45,336 questions and 28,408 images that necessitate reasoning about text for answering.
 We use the validation set, containing 5,000 images, for evaluation.
-4) ChartQA (Masry et al.
-2022) is a large-scale benchmark featuring 20,882 charts, with questions focusing on logical and visual reasoning.
-5) DocVQA (Mathew, Karatzas, and Jawahar 2021) concentrates on document image understanding, encompassing 50,000 questions and over 12,000 images.
+4) ChartQA [36] is a large-scale benchmark featuring 20,882 charts, with questions focusing on logical and visual reasoning.
+5) DocVQA [37] concentrates on document image understanding, encompassing 50,000 questions and over 12,000 images.
 We evaluate using the validation set, which includes 5,349 questions and 1,286 images.
-6) Seed2 Plus (Li et al.
-2024a) specifically designed for text-rich visual comprehension evaluation of MLLMs, includes 2.3K multiple-choice questions covering charts, maps, and webs.
-Figure 3: The visualization results of examples sampled from BLINK and RealWorldQA.
-Comparisons with State-of-the-art Models
-General Benchmarks.
-In Tab.
-2, we compare EE-MLLM with various state-of-the-art MLLMs (Liu et al.
-2024; Bai et al.
-2023; Ma et al.
-2024), across eight general benchmarks.
-These benchmarks evaluate the comprehensive ability of MLLMs, including understanding and perception, as well as the severity of illusion.
-These factors collectively reflect the generalization ability and applicability of MLLMs in real-world scenarios.
-EE-MLLM achieves comparable performance with state-of-the-art MLLMs in general benchmarks.
-Specifically, EE-MLLM achieves a score of 70.4 in MMBench and 1528.1 in MME, scores that are notably higer compared to LLaVA-v1.6, which also support high-resolution images input.
-This suggests that EE-MLLM possesses comprehensive perception and reasoning capabilities.
-What’s more, EE-MLLM also achieves promising performance on CCBench and SeedBench.
-Figure 4: Illustration of Inference Speed Ratio between EE-MLLM and LLaVA for input resolutions of $980\times 980$, considering numbers of generated tokens ranging from 2 to 256, for single and double image inputs.
-Fine-grained Benchmarks.
-In Tab.
-3, we perform evaluations on seven fine-grained benchmarks.
-These benchmarks necessitate impressive visual perception capability of MLLMs, as they require the exploration of fine-grained information within images to answer the questions.
-For traditional VQA benchmarks, such as TextVQA, ChartQA, and DocVQA, MLLM achieves a very impressive performance, 4.6 higher on TextVQA and 13.0 higher on ChartQA compared to LLaVA-v1.6, which also supports high resolution.
-In the OCRBench, specifically designed to assess the OCR capabilities of MLLM, EE-MLLM outperforms by 4.2 compared to LLaVA-v1.6.
-These results demonstrate that despite EE-MLLM significantly reducing the computational overhead associated with visual tokens, it effectively maintains the model’s fine-grained capabilities.
-Comparisons of Inference Speed
-Despite the significant reduction in FLOPs demonstrated by EE-MLLM, it is crucial to consider real-world scenarios where advanced techniques such as KVCache and Batch Inference are typically employed during deployment.
-Hence, conducting comparisons of inference speed with these techniques becomes necessary.
-We conduct the comparisons of inference speed on a single NVIDIA H800.
-The resolution of the input image is set to $980\times 980$, and the number of generated tokens varies from 2 to 256.
-We illustrate the speed ratio between EE-MLLM and LLaVA in Fig.
-4.
+6) Seed2 Plus [24] specifically designed for text-rich visual comprehension evaluation of MLLMs, includes 2.3K multiple-choice questions covering charts, maps, and webs.
+10 Comparison of Overall Inference Speed
+For MLLMs, the overall inference time consists of the prefilling time and the decode time.
+The prefilling time mainly involves processing visual inputs and textual instructions, while the decode time refers to the duration required to generate new tokens.
+In the main text, we have demonstrated EE-MLLM’s substantial improvements over LLaVA in reducing prefilling time.
+Here, we further compare the overall inference speed.
+In Fig.
+6, we present the speed ratio between EE-MLLM and LLaVA when generating varying numbers of tokens.
+We conduct the comparisons on a single NVIDIA H800.
+The resolution of the input image is set to 980×980, and the number of generated tokens varies from 2 to 256.
 Our findings show that when generating 8 tokens, EE-MLLM’s inference speed is three times that of LLaVA.
@@
 However, the advantage in inference speed diminishes after the first token.
-Specifically, with two input images, the inference speed of EE-MLLM is nearly four times that of LLaVA.
-This clearly indicates that EE-MLLM is significantly more efficient in multi-image input scenarios, including interleaved image-text dialogues and multimodal in-context learning.
-Stage EE-MLLM LLaVA
-Pre-training 32G 75G
-Fine-tuning 66G 69G
-Table 5: Memory usage of GPU during pre-training and fine-tuning stages.
-GPU Memory Overhead of EE-MLLM
-We assess the GPU memory overhead of EE-MLLM during both the pre-training and fine-tuning stages.
-Similarly, we compare EE-MLLM with LLaVA at a resolution of $336\times 336$.
-We perform experiments on 8$\times$×H800, with a global batch size of 256 for pre-training, 128 for fine-tuning.
-The comparison of memory usage is detailed in Tab 5.
-During the pre-training stage, EE-MLLM exhibits significant lower memory usage, at 32G, while the memory usage of LLaVA is 75G.
-During the fine-tuning stage, as the primary memory usage is focused on the trainable LLM, the memory usage advantage becomes less noticeable, with EE-MLLM consuming 66G and LLaVA-v1.6 utilizing 69G.
-Ablation Study
-Implementation Details
-Following LLaVA-v1.5 (Liu et al.
-2023a), we employ Vicuna-7b-v1.5 (Zheng et al.
-2023) as our foundational LLM.
-The training data is consistent with LLaVA-v1.5, encompassing both pre-training data and supervised fine-tuning data.
-Comparisons with LLaVA
-In Tab.
-4, we compare EE-MLLM with LLaVA under the same settings across both general and fine-grained benchmarks.
-For the resolution of $336\times 336$, we employ CLIP-ViT-L-14 as the vision encoder, ensuring complete alignment with LLaVA-v1.5.
-EE-MLLM demonstrates comparable performance to LLaVA on general benchmarks.
-In terms of fine-grained benchmarks, at $336\times 336$ resolution, EE-MLLM exhibits promising performance on AI2D and ChartQA but slightly underperforms LLaVA on OCRBench and TextVQA.
-The average score for the $336\times 336$ resolution is 47.1, which is 98% of LLaVA’s 48.1.
-For the resolution of $672\times 672$, we utilize SigLIP as the vision encoder.
-As shown in Tab.
-4, EE-MLLM obtains comparable results on AI2D and TextVQA, and the average score maintains 98% of LLaVA’s performance.
-V O FFN SeedB Hall TextVQA ChartQA DocVQA Avg.
-$336\times 336$
-65.3 24.4 44.8 14.4 18.4 33.5
-66.3 25.9 45.5 16.7 19.2 34.7
-65.4 25.3 45.4 15.6 19.1 34.2
-65.8 25.1 43.5 14.8 17.9 33.4
-66.1 25.7 46.1 15.8 18.6 34.5
-$672\times 672$
-66.1 25.2 56.8 19.7 28.6 39.3
-66.6 25.5 58.2 23.0 29.8 40.6
-66.1 24.3 58.2 21.6 30.1 40.1
-66.4 24.1 56.3 18.5 26.5 38.4
-66.9 27.6 58.8 24.0 31.8 41.8
-Table 6: Ablation of the aligner. Checkmarks indicate the weights employed in the aligner.
-Ablation of the Aligner
-In Tab.
-6, we evaluate the aligner variants with different mapping weights.
-We employ SigLIP initialized from Idefics2 (Laurençon et al.
-2024) to conduct experiments at different input resolutions.
-We remove the different weights from the aligner and train the model at a resolution of $336\times 336$.
-Checkmarks indicate the weights employed in the aligner.
-We have three findings: 1) Removing the entire aligner, as depicted in the first row, significantly reduces performance across multiple benchmarks.
-Specifically, the TextVQA score drops from 46.1 to 44.8, and the average score for the five benchmarks decreases from 34.5 to 33.5.
-This result highlights the aligner’s effectiveness in aligning visual features with the LLM feature space, allowing text tokens to capture crucial visual information and address questions through causal cross-attention modules.
-2) When ablating individual weights in the aligner, we find that maintaining the structure is more critical.
-The absence of $V$ or $O$ has a relatively minor impact on low-resolution input, even resulting in slightly higher average performance when $V$ is missing.
-However, when the FFN is missing, the aligner’s structure no longer resembles a transformer block, leading to a significant performance loss.
-3) We directly increase the input image resolution to $672\times 672$ without additional training and compare the variants with different aligner types.
-We observe that, at high-resolution input, the absence of $V$ or $O$ weights results in a substantial decline in fine-grained benchmarks such as TextVQA, ChartQA, and DocVQA.
-This finding demonstrates the importance of a complete aligner when applied to high-resolution images.
-Visualization
-We present four examples sampled from BLINK (Fu et al.
-2024) and RealWorldQA (XAI 2024) to assess the impact of the architectural changes in Fig 3.
-The first example demonstrates that EE-MLLM can perceive fine-grained visual context within the image, such as the color of traffic lights.
-The second and third examples highlight EE-MLLM’s ability to comprehend the position of objects.
-Specifically, EE-MLLM can accurately identify guitar’s position relative to the table and the dog’s position.
-The last example reveals that EE-MLLM can distinguish subtle differences in visual content.
-Conclusion
-In this paper, we revisit previous studies on multimodal large language models and categorize them into two groups: self-attention-based methods and cross-attention-based methods.
-The former is data-efficient but computationally inefficient, while the latter is compute-efficient but data-inefficient.
-To ensure data-efficient while maintaining compute-efficient, we propose the composite attention mechanism for EE-MLLM, which incorporates the composite attention module for computational efficiency and the aligner for data efficiency.
-We conduct comprehensive experiments on general benchmarks and fine-grained benchmarks and find that EE-MLLM achieves state-of-the-art performance on most benchmarks.
-We also assess the inference speed of EE-MLLM in real-world scenarios, and the results demonstrate that EE-MLLM holds a significant advantage in inference.
+Although the inference speed of EE-MLLM decreases as the number of generated tokens increases, it still maintains an advantage over LLaVA, especially under multi-image input scenarios.